OpenStudy (anonymous):

Graph and check to solve the system of linear equations.y = 2x – 4y = -1/2x + 1

OpenStudy (anonymous):

helpp

OpenStudy (anonymous):

i am going to assume that it is \[y=2x-4\\y--\frac{1}{2}x+1\] am i right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

rather \[y=2x-4\\y=-\frac{1}{2}x+1\]

OpenStudy (anonymous):

since \(y=y\) you can write \[2x-4=-\frac{1}{2}x+1\] and solve for \(x\)

OpenStudy (anonymous):

if the fraction annoys you, multiply by 2 and get \[4x-8=-x+2\] maybe that makes it easier

OpenStudy (anonymous):

huh? im trying to get points for x and y

OpenStudy (anonymous):

..... can you explain again please

OpenStudy (anonymous):

first find x then find y

OpenStudy (mathstudent55):

Instructions are: "Graph and check to solve the system of linear equations..."

OpenStudy (mathstudent55):

Let x = 0 and 1 in each equation, and find y. Fill in the tables below the equations. This will give you two (x, y) points for each equation. Then plot the two lines, and see where they intersect. that point is the solution. To check, test the point of intersection in both original equations to make sure it makes them true.